OpenStudy (anonymous):

Find the number of ways to listen to 4 different CDs from a selection of 15 CDs?

ganeshie8 (ganeshie8):

You can choose \(4\) CDs from a collection of \(15\) CDs in \(\binom{15}{4}\) ways. Each of these groups of \(4\) CDs can be listened in \(4!\) ways, so the total number of ways to listen \(4\) CDs from a collection of \(15\) CDs is \(\binom{15}{4}*4!\)

ganeshie8 (ganeshie8):

Alternatively, you can see that there are \(15\) choices for the first CD. After that, there are \(14\) choices for the second CD. After that, there are \(13\) choices for the third CD. After that, there are \(12\) choices for the fourth CD. So, the total number of ways to listen to \(4\) CDs from a collection of \(15\) CDs is simply \[15\times 14\times 13\times 12\]

OpenStudy (anonymous):

32,760?

OpenStudy (anonymous):

thank you

ganeshie8 (ganeshie8):

Looks good!

OpenStudy (usukidoll):

correct 15 x 14 x 13 x 12 = 32,760